### v2

# Element2Vec: Learning Chemical Element Representation from Text for Property Prediction

Abstract
Accurate property data for chemical elements is crucial for materials design and manufacturing, but many of them are difficult to measure directly due to equipment constraints. While traditional methods use the properties of other elements or related properties for prediction via numerical analyses, they often fail to model complex relationships. After all, not all characteristics can be represented as scalars. Recent efforts have been made to explore advanced AI tools such as language models for property estimation, but they still suffer from hallucinations and a lack of interpretability. In this paper, we investigate Element2Vecto effectively represent chemical elements from natural languages to support research in the natural sciences. Given the text parsed from Wikipedia pages, we use language models to generate both a single general-purpose embedding (Global) and a set of attribute-highlighted vectors (Local). Despite the complicated relationship across elements, the computational challenges also exist because of 1) the discrepancy in text distribution between common descriptions and specialized scientific texts, and 2) the extremely limited data, i.e., with only 118 known elements, data for specific properties is often highly sparse and incomplete. Thus, we also design a test-time training method based on self-attention to mitigate the prediction error caused by Vanilla regression clearly. We hope this work could pave the way for advancing AI-driven discovery in materials science.
keywords:
Large language model , Material science , Element encoding , Deep learning , Text mining , Data science
journal: ICLR 2026
[…]
[1] organization=Songshan Lake Materials Laboratory, city=Dongguan, state=Guangdong, country=China
[…]
[2] organization=International Academic Center of Complex Systems, Beijing Normal University, city=Zhuhai, state=Guangdong, country=China
[…]
[3] organization=Department of Applied Physics, The Hong Kong Polytechnic University, city=Hong Kong, country=China
[…]
[4] organization=Department of Computer Science, City University of Hong Kong (Dongguan), city=Dongguan, state=Guangdong, country=China
[…]
[5] organization=Department of Computer Science, City University of Hong Kong, city=Hong Kong, country=China
1 Introduction
Accurate property data of chemical elements, e.g., atomic radius and conductivity, are of paramount importance to materials design and manufacturing, e.g., barrier material design. Nevertheless, the direct measurement of specific properties for certain elements presents inherent challenges, e.g., “Lithium-Ion diffusion barrier” where measuring ionic conductivity of Lithium requires complicated synthesis for dense ceramic pellets and thin films while the results are highly sensitive to materials defects. As an alternative, numerical analysis that leverages other properties of the element, e.g., anion size or analogous properties of related elements, e.g., Ti, Ge, Zr, is employed, which may still require profound expert knowledge and fail to model intricate element-wise relationship comprehensively, resulting in limited effectiveness.
With the recent advances of deep learning, the graph neural networks have emerged as powerful tools to implicitly learn element correlation for property prediction [1, 2, 3, 4, 5, 6, 7], but may always require the data of properties used in input to be complete. In contrast, the data in materials science is of inherent and severe sparsity. Then, the large language models (LLMs) are finetuned for automatic chemical knowledge extraction from literature, followed by encoding in pre-defined formats [8, 9, 10], and direct material property prediction from chemical composition [11, 12]. Nevertheless, both approaches still rely on one-hot or handcrafted descriptors for elements, hindering the generalizability. Furthermore, issues like hallucination [13] and data distribution shift can compromise the reliability and utility of LLM further.
Figure 1: Encode Element Representation from Text based on inspiration from Image Representation. (Left) Conventional image representation can be either a global embedding extracted from the whole image, or a set of local embeddings extracted from a sub-region that focuses on several key components [14], such as tail and head. (Right) Analogously, given the tremendous amount of text in element description and its diverse application and aspects, we can also extract a single Global embedding from all text or following the definition in [15] to extract Local embeddings highlighting different element attributes like atomic, chemical, thermal, and optical.
In this work, we propose to employ LLM as an embedding extractor and construct a vector representation of chemical elements from free-form textual descriptions. For one thing, the materials are multifaceted entities with the description requiring a synthesis of physical, chemical, and structural attributes [16]. Then, the many properties, e.g., phase behavior such as gas and liquid, as well as the material design process, cannot be intuitively represented as mathematical descriptors. For another thing, the strong representation encoded with multifaceted characteristics may enable application of a simple algorithm for reliable and explainable prediction [17, 18]. Thus, for each element, by integrating diverse information into a unified, context-rich representation [19], we may enable broader downstream applications.
Specifically, we propose Element2Vecthat leverages the LLM to extract embeddings from text parsed from a Wikipedia webpage, and these embeddings can be directly compared in the same representation space. As illustrated in Fig. 1 (Left), motivated by the image encoding techniques that produce both a single embedding to describe entire image (globally) and a set of localized embeddings focusing on specific sub-regions, for each element, in addition to a Global embedding from the entire Wikipedia page to capture holistic information, we also design the streamline to extract a set of Local embeddings, where each is tailored to a specific attribute category, e.g., chemical and atomic, following the taxonomy defined in [15]. Without loss of generality, we assess properties framed as both classification and regression tasks. In particular, for regression under high data sparsity, we propose a test-time training method based on self-attention that recasts prediction as an imputation problem to enhance precision. With the contribution summarized as follows, we hope that our efforts could open new avenues for accelerating materials design [20]:
  • We investigate the representations of chemical elements from text by leveraging the capabilities of LLM. Moving beyond numerical data, we believe that integrating diverse information formats into a unified representation is essential for improving downstream tasks like property estimation.
  • We investigate Element2Vec, a training-free framework that generates and compares a Global embedding distilled from the entire text corpus, i.e., Wikipedia, and Local embeddings that highlights specific attributes.
  • Property estimation involves both problems in the form of classification and regression. For regression under high data sparsity, we propose a test-time training method based on self-attention and formulate the problem as imputation.
2 Related work
AI for materials discovery. Deep learning graph networks such as CGCNN [1] and MEGNet [3] have been applied for property prediction [2], while bond angles are further incorporated to improve predictions [21, 22]. Unsupervised natural language processing techniques have also uncovered chemical knowledge [23]. With the recent success of language models, MatSciBERT [24] is designed and tailored to materials text for information extraction, and LLMs like GPT-3 [25] are fine-tuned on chemistry tasks to predict molecular and materials properties competitively in low-data regimes [8, 9, 10].
Chemical elements embedding. Early work used physical properties or compositional data to project elements into vector spaces that reflect periodic trends. For instance, methods based on self-organizing maps [26] and Magpie features [27] successfully grouped elements by similarity, while unsupervised approaches like Atom2Vec [28] and SkipAtom [29] learned embeddings from compound structures that mirrored the periodicity of properties in an unsupervised manner. However, the dimensionality-reduction techniques [28, 30] primarily rediscover known relationships from existing data, lacking a mechanism for extrapolation beyond established knowledge. Nevertheless, all of the methods mentioned above depend on fixed or task-specific elemental encodings.
Despite prior explorations with word embeddings [18] and materials knowledge graphs showing that language captures periodic trends and structure–property relations [23, 31, 32, 33], the graph model still privileges geometric features but ignores language-informed semantics [34, 22, 35]. Following the great success of large language models, recent efforts emphasize text mining, KG construction, or text-guided generation [20, 36]. Very recent “universal atomic embeddings” are trained from crystal data rather than textual descriptions [19]. Still, the gap in constructing element-level embeddings from text still exists, which motivate our Element2Vec approach that learns broadly applicable element embeddings from text and evaluates them in material property prediction.
Figure 2: Element2Vec-Locals Embeddings Extraction. For each element, given the text parsed from the Wikipedia webpage, we use LLM to label each sentence about the attribute categories [15] independently and summarize the full text to shorten the words via the prompts $\mathbf{p}_{c}$ and $\mathbf{p}_{g}$ respectively. For each subset, it is concatenated with the summary and fed into $f_{emb}$ for extraction.
Large Language Models (LLM) have revolutionized text processing, enabling the flexible extraction of structured knowledge from unstructured text [37, 25]. Their capabilities in document segmentation, section identification, and intent classification allow for the precise capture of property-bearing statements from scientific literature [38, 39, 40]. Research in scientific NLP has developed domain-specialized LLMs to extract structured knowledge from text [41, 42, 43].
This has spurred the development of domain-specialized LLMs, like SciBERT [44, 39], which demonstrate superior performance in scientific information extraction. Techniques such as prompting further enable complex tasks like classification and rationale generation without task-specific model fine-tuning. In our pipeline, we leverage these advancements to process Wikipedia text: an LLM segments it into attribute-specific units and filters noisy sentences, thus generating clean, structured input for subsequent embedding and prediction tasks.
Text-embedding models convert textual descriptions into fixed-size vectors that preserve semantic relationships [45]. Early word embeddings like word2vec utilize word co-occurrence to discern word meanings and identify patterns [46]. Nevertheless, they encounter difficulties in comprehending extended sentences and acquiring detailed knowledge about specific entities. arora2017simple [47] proposed SIF as a simple yet strong baseline for sentence embeddings, and logeswaran2018efficient [48] introduced QuickThoughts with a contrastive context-prediction objective. However, many text-only pipelines can be trained upon a sufficient amount of data and do not necessarily utilize structured knowledge, e.g., Wikipedia2Vec [49], ERNIE [50], KEPLER [51]. Recent large-scale embedding models leverage LLMs for data generation or as backbones [52, 53]. In this work, we use Gemini Embedding, which is trained atop the Gemini LLM and explicitly designed to retain knowledge within fixed-size vectors [54].
3 Element2Vec: Chemical Elements Representation from Text
For each chemical element, we utilize the text parsed from its Wikipedia webpage to construct the representation and consider both Global and Local embeddings.
Element2Vec-Global. As an intuitive baseline, given the text corpus $\{\mathbf{s}_{i}\}_{i=1}^{N_{e}}$ of element $e$ that consists of $N_{e}$ sentence $\mathbf{s}_{i}$. The Global embedding is defined as a holistic representation of the entire corpus, which can be obtained as $f_{emb}(\{\mathbf{s}_{i}\}_{i=1}^{N_{e}})$ where $f_{emb}$ is implemented as Gemini embedding [54].
Despite the Global embedding’s aim for a comprehensive representation, a single embedding may struggle to explicitly represent all distinct attributes [15] covered in the content. Meanwhile, the order at the attribute level depends on the sentence order, as the original sequential order of sentences preserved in the webpages also varies, the order-sensitive nature of $f_{emb}$ architecture can then introduce noise into the embeddings.
Element2Vec-Locals is thus proposed to highlight attributes in the extracted embedding without losing the global context. For efficiency purposes and without LLM retraining, we extract from a combination of a global summary and an attribute-specific subset, i.e., a filtered collection of sentences, as illustrated in Fig. 2.
In detail, following the concepts in [15], we define eight attribute categories, i.e., “Mechanical (Physical)”, “Optical”, “Electrical and Magnetic”, “Thermal”, “Chemical”, “Atomic and Radiational”, “Applications”, and “Abundance”. Then, we classify each sentence $\mathbf{s}_{i}$ using a prompt $\mathbf{p}_{c}$ that instructs the LLM to assign it to an attribute, i.e., $\text{LLM}(concat(\mathbf{p}_{c},\mathbf{s}_{i}))$. Sentences tagged with the same attribute are automatically grouped into a distinct subset. Finally, by concatenating with a concise global context, $\text{LLM}(concat(\mathbf{p}_{s},\{\mathbf{s}_{i}\}_{i=1}^{N_{e}}))$, summarized via LLM by the prompt $\mathbf{p}$, we extract the corresponding Gemini embedding [54]. The justification for including both components is provided by the ablation studies in Sec. 4.1.
Figure 3: t-SNE visualization of element2vec representations. (a) Embeddings obtained by sending the entire Wikipedia page with a prompt. (b) Local embeddings obtained from 8 attributes without summary. (c) Local embedding, one point per element obtained from the attribute plus its summary. (d) Aggregated embeddings from (c), 8 attribute-level embeddings are combined into a single point per element. (e) Global embedding: one point per element obtained directly. (f) Periodic table with the same color legend as in (a–e).
4 Property Prediction by Element2Vec
We first evaluate the characteristics of the Global and Local embeddings via element classification Sec. 4.1 and  Sec. 4.2. Subsequently, we apply these embeddings to property estimation, framing it as a regression problem and introducing and demonstrating the efficacy of the test-time training approach Sec. 4.3 and Sec. 4.4 even under conditions of high data sparsity.
4.1 Classification of chemical elements
In this paper, we define family as a group of elements in the periodic table that share similar chemical properties, electronic configurations, and reactivity patterns. Specifically, families are defined as the vertical columns (element groups), e.g., Alkali metals (Group 1: Li, Na, K, …). Family classification is the process of categorizing elements into these families.
We use t-SNE to compare Element2Vec-Global and Element2Vec-Locals representations under identical projection hyperparameters and a fixed random seed (Fig. 3). Fig. 3(a) feeds the entire Wikipedia page to the Gemini embedding model with an added instruction prefix (“prompt”) intended to induce attribute-aware encoding. Fig. 3(b) uses an LLM to segment the page into eight predefined attributes and embeds each attribute text without any summary. Fig. 3(c) applies our Element2Vec-Locals method, in which each attribute text is concatenated with the element-level summary before embedding; the inset figure in Fig. 3(c) highlights two elements to show that the eight attribute points of the same element form a compact local cluster rather than collapsing to a single point. Fig. 3(d) aggregates the eight attribute embeddings from Fig. 3(c) into one vector per element by concatenation (details in Sec . 3), whereas Fig. 3(e) shows the Element2Vec-Global pipeline where one vector is obtained directly from the full page. Fig. 3(f) provides the periodic table with the same color legend; colors denote chemical families, and markers distinguish lanthanides (triangles) and actinides (squares).
Three observations follow. First, in Fig. 3(a), the attribute-level points of a given element nearly coincide, indicating that adding a prompt to the full-page input does not induce meaningful attribute selectivity; the representations behave almost as if the same text were embedded repeatedly, which means that our prompt did not work in Gemini embedding. Second, in Fig. 3(b), the eight attributes per element scatter broadly and do not form identifiable per-element clusters, suggesting that removing the shared global context makes attribute snippets too heterogeneous to anchor to their source element. Third, Fig. 3(c) yields a markedly more coherent geometry: for nearly all elements, the eight attribute points cluster together while retaining small intra-cluster spread, consistent with the design goal of preserving both shared element identity (via the summary) and attribute-specific variation (via the segmented text).
At the element level, Fig. 3(d) (Element2Vec-Locals) aligns more closely with the periodic families than Fig. 3(e) (Element2Vec-Global). Families such as metalloids, alkali metals, nonmetals, and alkaline earth metals form tighter, cleaner groups in Fig. 3(d), and transition metals—including lanthanides and actinides marked by different symbols—also exhibit improved compactness. Because all panels share identical t-SNE settings, these differences reflect the representations themselves. In summary, summary-augmented Element2Vec-Locals embeddings produce attribute-level micro-clusters anchored on their parent elements and yield Element2Vec-Locals that respect periodic families more faithfully than Element2Vec-Global.
Figure 4: Entropy distributions for family classification. Kernel density estimation (KDE) of prediction entropy for the 10-way periodic-family classifier under three embeddings.
4.2 Entropy analysis with summary-augmented Element2Vec-Locals embeddings
Based on Fig. 3(d) and Fig. 3(e), where Element2Vec-Locals yielded tighter intra-element clusters and clearer family-level separation than Ement2Vec-Global, we now move from qualitative visualization to a quantitative assessment of prediction decisiveness.
Encoding an entire element page into a single vector can dilute attribute-specific signals, whereas encoding only attribute snippets may lose the global context that relates an element’s properties. We therefore adopt an Ement2Vec-Locals scheme augmented with an element-level summary: each Wikipedia page is segmented into eight predefined attributes; a concise summary of the full page is produced. For each attribute, the attribute’s text is concatenated with the summary (front or end), encoded with the Gemini embedding model, and used in downstream tasks. We evaluate three datasets: Element2Vec-Global (global), Ement2Vec-Locals with summary prepended (local, front), and Ement2Vec-Locals with summary appended (local, end).
To quantify predictive uncertainty for the 10-way family classifier, we compute the Shannon entropy of the softmax posterior for each test element $x$,
$H(x)=-\sum_{k=1}^{K}p_{k}(x)\,\log_{2}\!\big(p_{k}(x)+10^{-9}\big),$ (1)
where $K=10$ and $p_{k}(x)$ is the predicted probability of class $k$; lower $H(x)$ indicate sharper, less ambiguous posteriors. The constant $10^{-9}$ is to ensure $\log_{2}0$ is never evaluated when $p_{k}(x)=0$.
The resulting distributions (Fig. 4) show that both local, front and local, end shift entropy toward lower values compared with global, indicating sharper posteriors and reduced ambiguity. Between the two summary placements, local, end consistently exhibits a slightly larger mass at lower entropies than local, front. A possible interpretation is that appending the summary preserves attribute-specific leading context while still injecting a shared global prior, yielding marginally more decisive predictions. Overall, augmenting attribute segments with a brief summary reduces uncertainty relative to a single global embedding, and placing the summary at the end offers a modest additional gain.
4.3 Test-Time Training for Robust Property Value Imputation
Figure 5: Pipeline comparison between (a) test-time training and (b) standard inductive training.
Figure 6: Evaluation of van der Waals radius prediction via local embeddings. (a) Schematic of interatomic interaction energy as a function of internuclear separation, showing the definition of the van der Waals radius ($R_{vdW}$). (b) Root mean square error (RMSE) as a function of embedding dimensions $N$ for different local embeddings of elemental attributes, with atomic and radiational features (ARF) highlighted in red. (c) RMSE versus data missing ratio under different ways of constructing the local embeddings.
Different from the family classification problem formulated above where the labels of all elements are given, a key objective for property prediction is to estimate the property values for certain elements which are unknown or hard to measure.
As an intuitive baseline and shown in Fig. 5 (b), the inductive training pipeline can be considered, where a predictor $f_{p}$, with typical architecture as linear layer or MLP, is first trained on elements $\mathcal{E}_{k}$ with known property value and directly tested on the other elements $\mathcal{E}_{uk}$ whose property values are missed, where $\mathcal{E}_{k}$ and $\mathcal{E}_{uk}$ stands for the set of embeddings for the corresponding element set, i.e., $|\mathcal{E}_{k}\cup\mathcal{E}_{uk}|=118$ and $\mathcal{E}_{k}\cap\mathcal{E}_{uk}=\emptyset{}$. Then, the core assumption is that the model will generalize from the training elements to unseen test elements, which could be nevertheless sub-optimal when the training data is limited. After all, for certain properties, the property data is highly sparse where the values are available for less-than-20% of the total 118 elements comprising the periodic table.
Motivated by the fact that the properties of an element can be influenced by those of others, it is crucial to model inter-element correlations for accurate prediction. Consequently, we propose to employ a network $f_{sa}$ based on self-attention architecture [55] and design a corresponding test-time training pipeline [56] as shown in Fig. 5(a) as the training of network parameters and the property prediction of elements in $\mathcal{E}_{uk}$ can be done simultaneously.
In detail, we formulate the $\mathcal{E}_{k}\cup\mathcal{E}_{uk}$ as input sequence and fed into the self-attention layer, where the corresponding output of each input embedding is projected via a shared linear layer and used as the predicted property value. Then, for the network supervision, the objective is to only minimize the loss, e.g., mean squared error [1] of elements in $\mathcal{E}_{k}$ and no supervision will be applied for elements in $\mathcal{E}_{uk}$.
In this way, different from the conventional inductive training pipeline, where the gradients used for predictor parameter update is only derived from elements $\mathcal{E}_{k}$ with known property values, the gradients to update the self-attention parameters are also calculated from elements $\mathcal{E}_{uk}$ with unknown property values, which essentially training the network parameters on $\mathcal{E}_{uk}$ in an unsupervised manner and convert the problem of prediction on $\mathcal{E}_{uk}$ to value imputation. Consequently, the model can be directly trained from scratch and then use the model output at the last step as the final prediction, mitigating the performance drop due to scarce data and the distribution shift between $\mathcal{E}_{k}$ and $\mathcal{E}_{uk}$.
Figure 7: Model performance comparison against an increasingly incomplete dataset. (a) For global embedding, the RMSE vs. missing ratio obtained from traditional training and test-time training. (b) RMSE vs. missing ratio predicted via linear regression for global, local-concatenated, and local-ARF embeddings (with 0.05 summary ratio put at the front).
Figure 8: The shared feature dependencies across different material properties. (a) Normalized RMSE curves for selected materials properties, along with the corresponding saturation dimensions required to reach convergence. (b) Clustered heat map of feature-overlap similarity among the 14 selected materials properties, which is measured by the size of the intersection between their Top-$k$ selected dimensions.
During implementation, as the true property values of elements in the set $\mathcal{E}_{uk}$ are naturally unknown and cannot be used to quantitatively evaluate the model performance, we instead simulate a scenario by helding out certain elements via random selection as a validation set from $\mathcal{E}_{k}$. Specifically, we partition $\mathcal{E}_{k}$ into a training set $\mathcal{E}_{tr}$ and a test set $\mathcal{E}_{te}$, such that $\mathcal{E}_{k}=\mathcal{E}_{tr}\cup\mathcal{E}_{te}$ and $\mathcal{E}_{tr}\cap\mathcal{E}_{te}=\emptyset{}$. The ratio $|\mathcal{E}_{te}|\ |\mathcal{E}_{k}|$ is defined as missing rate, reflecting the difficulty of the property estimation.
For material research, since the atomic number, i.e., the indexes of the elements in the periodic table, is a critical factor in property prediction, the position encoding [55] should have been incorporated into our algorithm as inputs. However, this is unnecessary from our observation and we think the reason is because that the embedding extracted from the source text already encapsulate the corresponding information, thus rendering explicit encoding redundant.
4.4 prediction of atomic properties
Following Section 4.2, to avoid confounding representation changes, all subsequent regression experiments therefore use the summary-prepended setting (local, front). Fig. 6 turns to a representative scalar property—the van der Waals (vdW) radius—both to motivate the task and to test whether text-derived embeddings carry element-level size information.
Fig. 6(a) shows the interaction–distance curve and the $R_{\mathrm{vdW}}$ relation used here, underscoring why $R_{\mathrm{vdW}}$ matter: they enter noncovalent contact a steric accessibility, packing, and force-field parameterization across chemistry, materials, and biomolecular modeling [57, 58, 59, 60]. At the same time, $R_{\mathrm{vdW}}$ is difficult to determine and not uniquely defined; estimates depend on environment and methodology, often inferred from contact distributions, dispersion models, or polarizability trends rather than measured directly [61, 62, 63]. This importance–difficulty trade-off motivates evaluating whether our embeddings can predict element-level $R_{\mathrm{vdW}}$.
In Fig. 6(b), we conduct an experiment to validate whether a specific property aligns with a specific attribute ($R_{\mathrm{vdW}}$-ARF). We first remove elements with any missing attribute text or missing target value, yielding 96 elements. For each of the eight local, front attribute embeddings, we run 10-fold cross-validation and report RMSE on held-out folds. To study feature sufficiency, in every fold, a linear regressor is trained after ranking dimensions on the training split and selecting the top $N$ features ($N\in[100,768]$; selection is performed within the training fold to avoid leakage), and a full-feature linear baseline is also fitted on the same split. As $N$ increases, RMSE decreases and stabilizes; among attributes, Atomic and radiational features attain the lowest error (red curve), consistent with the expectation that $R_{\mathrm{vdW}}$ correlates with atomic size and electronic extent captured by this attribute.
In Fig. 6(c), we vary the summary-to-page ratio in local, front ($0.05,0.10,0.20$) and probe robustness under scarce data by sweeping the missing ratio (test fraction) up to $80\%$. For each level, we randomize train/test splits five times and average results. Because RMSE can fluctuate with feature count, we aggregate performance using the best-tail average RMSE: identify $N_{\text{best}}$ with the minimum RMSE and average RMSE over $N\in[N_{\text{best}},N_{\max}]$. As shown in Fig. 6(c), when the missing ratio is below $50\%$, the $0.05$ summary consistently performs best, suggesting that a shorter summary preserves attribute-specific signal while supplying a light global prior. Between $50\%$ and $80\%$, the $0.20$ summary becomes superior, indicating that stronger shared context helps stabilize regression when training evidence is limited and attribute coverage is uneven. Overall, summary-augmented Element2Vec-Locals embeddings support $R_{\mathrm{vdW}}$ prediction, with shorter summaries favored in data-richer regimes and longer summaries acting as a useful prior as data become scarce.
Followingly, given the limited dataset available for inferring atomic-level properties, we keep using the global embedding and then compare the performance of different predictive models in Fig. 7.
First of all, we observe that test-time training consistently performs better than the traditional training approaches (both linear regression(LR) and MLP), particularly as the missing ratio increases. When comparing LR and MLP, when the labeled data is pretty limited, aggressively increasing the size of the network will actually make the prediction even worse. In contrast, our test-time training algorithm can adapt the model effectively to mitigate the impact of incomplete inputs, leading to substantially lower RMSE values across all missing ratios, especially at the higher ones.
In Fig. 7(b), we show the trends of the influence of different embedding strategies. Fig. 6(c) has already shown that 0.05 summary putting at the front brings the lowest error among the Element2Vec-Locals. Therefore, we decide to show these Element2Vec-Locals results here for comparison against the Element2Vec-Global. Interestingly, the global embedding generally exhibits the lowest error across most missing ratios, indicating that Element2Vec-Global are inherently more accurate than localized representations for property predictions. These insights point to the combined potential of Element2Vec-Global with test-time training as a promising direction for building reliable models under real-world conditions where data incompleteness is unavoidable. The selected representative elements’ predictions of $R_{\mathrm{vdW}}$ via Element2Vec-Global with 95% confidence interval are presented in Table 3 in the Appendix.
Finally, to establish ranking among different global embedding dimensions for a more generalized material property prediction, we perform linear regression with a feature-budget analysis. For a given target material property $p$, we first fit the linear model to all 768 dimensions in the global embedding and rank dimensions by the absolute value of standardized coefficients $w_{j}=|\beta_{p,j}|$. For each budget $k\in\{100,101,\dots,768\}$, the linear model is retrained from scratch using only the top-$k$ ranked dimensions $\{r_{p}(1),\dots,r_{p}(k)\}$ (Fig. 8(a)). Generalization is assessed via 10-fold cross-validation as described above.
In Fig. 8(b), we select 14 representative material properties and visualize the overlap of their Top-$k$ dimensions using a clustered heatmap. Interestingly, boiling point and melting point share the highest degree of overlap, reflecting their close thermodynamic relationship, whereas their intersection with other thermal properties (e.g., thermal conductivity) is minimal. This observation suggests that while some properties naturally cluster due to shared underlying physical principles, others remain more distinct. More broadly, this finding highlights how the learned embeddings capture both intrinsic scientific connections and the way such relationships are expressed in human language, thereby bridging the gap between physical attributes and linguistic representation.
5 Conclusion
In this paper, we have proposed Element2Vec, an LLM-based framework for embedding chemical elements that turns human-experienced knowledge into structured materials information. By leveraging the Wikipedia page as a resource, we construct both Global and Local attribute-specific embeddings that capture the correlations among elements. Our experiments demonstrate that these embeddings can recover periodic trends and improve element classification. We also showcase the stability imputation of missing values through test-time training via an attention network. Beyond performance gains, the framework integrates large language models with domain-specific structure to uncover latent knowledge and provide transferable, interpretable representations. Future directions will be addressed to build foundation model representations for more generalizable tasks in materials science, such as the prediction of material composition and the synthesis process. The work paves the way for new materials discovery by turning unstructured scientific knowledge into physics-informed representations.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (52471178).
Appendix A Learning elemental properties by multi-layer perceptron
The multilayer perceptron (MLP) used in this study has a three-layer architecture consisting of an input layer, a single hidden layer with 100 neurons using ReLU activation, and an output layer. It is configured for efficient training using the Adam optimizer with very weak L2 regularization (alpha=0.0001). The model employs a maximum of 500 iterations and has early stopping enabled to prevent overfitting by halting training if the validation score stops improving. Given that the abbreviations of attribute names are summarized in Table 1, we summarize the statistics of word count for each local attributes in the Table 2. The detailed performance of all local embedding datasets are summarized in Fig. 1.
Table 1: Attribute names and abbreviations
Full name Abbr.
Mechanical properties (Attr.1) MECH
Optical properties (Attr.2) OPT
Electrical & Magnetic properties (Attr.3) EM
Thermal properties (Attr.4) THERM
Chemical properties (Attr.5) CHEM
Atomic & radiational features (Attr.6) ARF
Applications (Attr.7) APPL
Abundance (Attr.8) ABND
Table 2: Word count of Element2Vec-Locals dataset with different summary ratio(S/R)
S/R Elem. Attr.1 Attr.2 Attr.3 Attr.4 Attr.5 Attr.6 Attr.7 Attr.8 Sum.
0.01 He 1234 212 540 2202 2081 731 2741 3112 442
C 1463 666 766 1535 7619 5652 4168 5677 291
Ca 414 37 185 363 3314 1837 1524 1569 244
Ar 182 656 987 579 4155 1823 3537 1508 162
Br 260 520 801 1475 9146 843 4411 1858 345
Au 1622 1217 1642 172 4782 1009 6346 6019 567
Nb 1294 617 2254 711 3856 582 2764 1086 256
Sm 613 1764 1960 2565 4750 3067 3468 1583 216
0.05 He 2519 901 758 1950 2644 2551 3043 5736 1119
C 1790 675 721 1452 6642 3761 4302 5691 1821
Ca 616 34 652 257 6012 3498 4215 709 1408
Ar 59 629 464 404 2655 1357 1560 644 646
Br 114 730 381 1664 9318 585 3561 1330 1730
Au 1425 1239 1897 273 $50$ 1885 6014 5381 2835
Nb 1245 1016 1957 702 3589 1107 5424 718 822
Sm 672 1853 2768 1299 4883 3682 3482 1927 1325
0.1 He 1996 800 1065 3663 3196 5414 5150 5838 2027
C 2658 746 714 1365 7653 6621 2518 5445 2620
Ca 793 34 620 478 5828 2784 3622 1795 2817
Ar 59 568 424 432 2674 1585 1467 486 1143
Br 244 647 703 3237 10074 957 4476 1525 3336
Au 1333 1061 1582 49 5752 1090 10126 6468 4486
Nb 1209 527 2225 529 4204 449 4761 888 1504
Sm 613 1545 2066 1554 4624 4690 3826 2366 2650
0.2 He 1608 470 802 3083 2249 1971 1872 3565 3383
C 1801 580 624 1182 7590 5409 2836 6681 2444
Ca 553 87 510 342 7695 3992 3952 1738 4369
Ar 251 471 403 534 2136 1735 2499 633 1380
Br 629 420 642 2417 11770 761 3657 1281 5104
Au 1950 1238 2060 278 4604 1060 5453 5570 4110
Nb 1051 814 2273 629 3732 583 3908 1063 3040
Sm 487 1451 2527 1053 5516 4218 3775 2107 2306
Fig. 1: (a) The LR result of all Element2Vec-Locals embedding datasets with full attributes vector. (B) The LR result of all Element2Vec-Locals embedding datasets with only ARF attribute vector.
Table 3: $R_{\mathrm{vdW}}$: ground truth and predicted values with 95% confidence interval. The values in the table are in Å; $1~\text{\AA }=10^{-10}\,\text{m}$.
Element True Value Predicted Value
He 1.40 1.42 $\pm$ 0.11
C 1.70 1.65 $\pm$ 0.12
Ca 2.31 2.27 $\pm$ 0.06
Ar 1.88 1.78 $\pm$ 0.10
Br 1.85 1.88 $\pm$ 0.08
Au 2.14 2.11 $\pm$ 0.06
Nb 2.18 2.31 $\pm$ 0.04
Sm 2.36 2.38 $\pm$ 0.06